Home CMS Production Clean room MedAustron HEPHY testbeams old
IrradXray-CERN-2022 testbeam-RD50-CERN-Oct2022 testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  HEPHY logbook for testbeam at SPS October 2012 and Irradiation Mol  Not logged in ELOG logo
Message ID: 35     Entry time: Wed Oct 17 21:30:39 2012
Author: Markus Friedl 
Project: MicronOrigami_gross 
Subject: cal 
Run Number: cal008 
Events: 83800 
StartTime: 17.10.2012 17:15:13 
EndTime: 17.10.2012 18:08:11 
Data: Good 

Internal calibration for "stealth parasitic" mode Belle stack; taken without beam


Belle Stack 4: 2 DSSD -z, Origami #3, #1, wedge new #2
HV = +-50V, Ibias=80A
Position: approx. centered
no beam


Total bias current history:


At the beginning, the system was already running for a few hours and the APV25 chips were initialized before, so thermal equilibrium was reached.

The actual calibration run happened during the first ~250 measurement points (~1 hour). It seems like the APVs draw a bit more current during internal calibration. The overall decay is supposedly due to a temperature decrease towards night.


Now tried the same thing with a short software run - same effect:


Need to investigate supply currents while data taking - they should actually go up a bit - and their dependency on the trigger rate.


ELOG V3.1.4-395e101